# vAttention: Dynamic Memory Management for Serving LLMs without PagedAttention

### v1

@@ 6.2. Hiding memory allocation latency
The serving framework invokes the step API in every iteration. The latency of step depends on how many new pages need to be mapped in the virtual tensors of KV-cache. Consider, for example, that the KV-cache of one request needs to be extended for Yi-34B which has 60 layers. This requires 120 calls to vMemMap each of which takes about 9 microseconds. Therefore, growing the KV-cache of one request by one page will add about 1 millisecond latency to the corresponding iteration and would grow proportional to amount of physical memory that needs to be mapped. We propose the following optimizations to hide the latency of allocation:
6.2.1. Overlapping memory allocation with compute
@@
6.2.2. Deferred reclamation + eager allocation
We observe that allocating physical memory for a new request can be avoided in many cases. Consider that a request R1 completed in iteration i and a new request R2 joins the running batch in iteration i+1. To avoid allocating new pages to R2 from scratch, vAttention simply defers the reclamation of R1’s pages and assigns R1’s reqId to R2. This way, R2 uses the same tensors for its KV-cache that R1 was using which are already backed by physical pages. Therefore, new pages for R2 are required only if its context length is bigger than that of R1.
We further optimize memory allocation by proactively mapping physical pages before they are needed. We do so by using one of the inactive reqId’s KV-cache. When a new request arrives, we can allocate this reqId without mapping any physical pages. We then select a new reqId that would be allocated next and map physical pages for it. In most cases, these eager optimizations obviate the need to allocate new physical pages even for the prefill phase of new requests. Finally, we trigger memory reclamation only when the number of physical memory pages cached in vAttention falls below a certain threshold (e.g., less than 10% of GPU memory). We delegate both deferred reclamation and eager allocation to the background thread that the step API spawns.
@@ 6.2.2. Deferred reclamation + eager allocation
7. Evaluation
@@
  • How does vAttention perform for prefill and decode phases in LLM inference? What are the portability and performance advantages of vAttention.
  • How efficiently can vAttention allocate GPU memory for LLM serving workloads, and how effectively can it deal with KV-cache fragmentation?
@@
Models and Hardware: We evaluate three models Yi-6B, Llama-3-8B and Yi-34B, using a single NVIDIA A100 GPU for Yi-6B, and two NVLink-connected A100 GPUs for Llama-3-8B and Yi-34B (see Table 5). Each GPU has 80GB physical memory. We use tensor-parallelism degree of two (TP-2) for both Llama-3-8B and Yi-34B. All three models use GQA which is the most commonly used attention mechanism in recent LLMs.
@@
Evaluation methodology: The computation and memory allocation pattern of the prefill and decode phases is substantially different. Attention kernels used for these two phases are also different and hence we evaluate them separately. The prefill phase requires one time memory allocation potentially spanning multiple pages. In comparison, the decode phase requires incremental memory allocation over the lifetime of a request. We measure the throughput of these phases in terms of tokens processed (or generated) per second.
Figure 8. Prefill tokens processed per second with chunking (chunk size=2048). Using vanilla prefill kernels, vAttention outperforms the paged counterpart of both FlashAttention and FlashInfer. Throughput for longer context lengths drops sharply due to quadratic complexity of prefill attention.
7.1. Portability and Performance for Prefills
To evaluate the prefill phase, we focus on the attention kernels provided by FlashAttention v2.5.6 (Dao, 2023; fag, 2022) and FlashInfer v0.0.3 (fig, 2023; Ye et al., 2024). We do not include vLLM in these experiments because it does not have a prefill kernel of its own but instead uses FlashAttention kernel. We also could not evaluate Yi-34B because FlashInfer kernels do not support Yi-34B’s KV group size of 7 (fip, 2024).
FlashInfer is a library that recently introduced a set of attention kernels optimized for different scenarios e.g., for chunked-prefills – an optimization proposed in Sarathi (Agrawal et al., 2023) and later adopted in various systems (Agrawal et al., 2024; Holmes et al., 2024; Hu et al., 2024). Sarathi splits the input tokens of a prompt into multiple smaller chunks and schedules one chunk at a time, enabling a serving system to add new requests in a batch without pausing ongoing decodes. This helps improve throughput without increasing latency (Agrawal et al., 2024). Both FlashAttention and FlashInfer provide kernels to compute the attention scores of chunked-prefills with and without PagedAttention. We integrate them into vLLM and using chunk size of 2048 tokens, measure time-to-first-token (TTFT) for the following configurations:
FA_Paged: Uses flash_attn_with_kv_cache kernel API of FlashAttention.
FI_Paged: Uses FlashInfer’s PagedAttention kernel, representing state-of-the-art PagedAttention-based kernel for the prefill phase.
FA_vAttention: Uses FlashAttention’s vanilla prefill kernel via the flash_attn_func API.
FI_vAttention: Uses FlashInfer’s vanilla prefill kernel via the single_prefill_with_kv_cache API.
Model Context Length FlashAttention FlashInfer
Paged vAttention Paged vAttention
Yi-6B 16K 2.44 1.38 1.51 1.34
32K 8.07 3.50 3.97 3.4
64K 30.46 9.98 11.85 9.66
128K 119 32.10 39.17 30.9
192K 266 67.81 81.87 65.03
@@ 7.1. Portability and Performance for Prefills
32K 5.19 2.59 3.06 2.37
Llama-3 64K 18.65 6.84 8.94 6.50
-8B 128K 71.16 20.78 28.98 20.29
192K 157.33 42.31 59.91 41.35
Table 6. Time-to-first-token (TTFT) for a single prompt of varying context lengths.
(a) Decode throughput (measured as the number of output tokens generated per second) with varying batch size.
(b) Mean GPU time spent in computing attention with varying batch sizes.
Figure 9. Decode throughput with varying batch sizes using context length 16K for each request (FA: FlashAttention, bs: block size). We evaluate vLLM and FlashAttention with two different block sizes: 16 and 128. vLLM performs best with block size 16 because its attention kernel is more efficient with smaller block sizes. FlashAttention’s GPU kernel is up to 2.85×2.85$\times$2.85 × faster than the best version of vLLM’s kernel (Yi-6B, 16*16K). However, smaller blocks add CPU overhead e.g., FlashAttention with block size 16 is worse than with block size 128. vAttention provides similar gains that the best version of FlashAttention provides over vLLM, but without user-level physical memory management and without a PagedAttention kernel.
Both vAttention configurations therefore use kernels that support chunked-prefills over a virtually contiguous KV-cache. We add dynamic memory allocation support to them without having to modify their code.
Figure 8 shows the prefill throughput of the four configurations for Yi-6B and Llama-3-8B. In all cases, vAttention delivers consistently higher throughput, outperforming FA_Paged by 1.603.92×1.60-3.92$\times$1.60 - 3.92 × and FI_Paged by 1.031.45×1.03-1.45$\times$1.03 - 1.45 ×. For the same experiment, Table 6 shows the TTFT with different context lengths. Since TTFT directly depends on prefill throughput, compared to using vanilla kernels with vAttention, FA_Paged and FI_Paged increase TTFT by up to 3.92×3.92$\times$3.92 × (Yi-6B, context length 192K) and 1.45×1.45$\times$1.45 × (Llama-3-8B, context length 192K), respectively.
The source of vAttention’s performance gain is twofold in these scenarios. First, the vanilla kernel is faster than the paged kernel in both FlashAttention and FlashInfer. While FlashAttention’s paged kernel is not optimized for prefills (it is optimized for decodes), FlashInfer’s paged kernel is specifically designed to support chunked-prefills. However, the paged kernel is slower than the vanilla kernel as discussed in §3.3. This example illustrates the complexities of transferring performance-critical optimizations between different implementations – even when the implementations are written by the same team. The second source of improvement is vAttention’s less CPU overhead. For example, appending a new K or V tensor to the KV-cache requires a single tensor copy operation in vAttention, whereas in a paged implementation, it requires appending one block at a time. Further, FlashInfer involves creation and deletion of a few objects for its compressed Block-Tables in every iteration. vAttention avoids such overheads because it maintains KV-cache’s virtual contiguity and therefore does not require a Block-Table.
7.2. Portability and Performance for Decodes
To evaluate decode performance, we focus on long-context scenarios (16K) because the latency of attention kernel becomes significant only for long contexts [^4] For short contexts, the computation time of the feed-forward-network dominates inference latency (Agrawal et al., 2024). We evaluate the following configurations:
vLLM: We use vLLM v0.2.7 as the primary baseline. vLLM pioneered PagedAttention and uses a custom paged kernel for decodes, derived from FasterTransformer (ftk, [n. d.]).
FA_Paged: For the second baseline, we integrate the FlashAttention kernel into vLLM’s serving stack. This represents a state-of-the-art PagedAttention kernel that includes optimizations such as sequence parallelism and in-place copy of new key and value vectors into the KV-cache. We evaluate the paged kernels of vLLM and FlashAttention with two different block sizes – 16 and 128 – to capture the effect of block size on performance.
FA_vAttention: For vAttention, we integrated the vanilla kernel of FlashAttention into vLLM’s serving stack. The kernel works with a virtually contiguous KV-cache to which we dynamically allocate physical memory using 2MB pages.
@@ 7.2. Portability and Performance for Decodes
9(a) shows the decode throughput of Yi-6B, Llama-3-8B and Yi-34B with varying batch sizes wherein the initial context length of each request is 16K tokens and we generate $256$ tokens for each request. We compute decode throughput based on the mean latency of 256 decode iterations. We summarize the key takeaways below.
First, vAttention outperforms vLLM (both block sizes) and FA_Paged (block size 16), while roughly matching the best configuration of FA_Paged (block size 128). The maximum improvement over vLLM is 1.97×1.97$\times$1.97 × for Yi-6B, 1.3×1.3$\times$1.3 × for Llama-3-8B and 1.6×1.6$\times$1.6 × for Yi-34B. The relative gains over vLLM also increase as the batch size grows. For example, the gain increases from about $\times1$ to 1.97×1.97$\times$1.97 × as batch size increases from 1 to 8 for Yi-6B. This is because the latency of attention computation grows proportional to the total number of tokens in the batch (see 9(b)) whereas the cost of linear operators remains roughly the same (Agrawal et al., 2023; Patel et al., 2023; Agrawal et al., 2024). Therefore, the contribution of attention kernel in the overall latency – and subsequently gain with a more efficient kernel – increases with the batch size. While FA_Paged (block size 128) provides similar gains as vAttention, note that FA_Paged requires a new implementation of the GPU kernel whereas vAttention simply leverages the vanilla kernel of FlashAttention.
Second, 9(b) confirms that performance difference between vLLM and FA_Paged/vAttention is indeed due to the attention kernels. In the worst case, the latency of vLLM’s best PagedAttention kernel (block size 16) is up to 2.85×2.85$\times$2.85 × higher for Yi-6B, up to 1.45×1.45$\times$1.45 × for Llama-3-8B and up to 2.62×2.62$\times$2.62 × for Yi-34B than the FlashAttention kernel.
Finally, throughput can be sensitive to block size even when memory capacity is not a constraint. For example, as discussed in §3.3, vLLM’s attention kernel has a significantly higher latency with block size 128 than with block size 16 (also see 9(b)). In the worst case, block size 128 degrades vLLM’s throughput by 36%. While block size has a smaller impact on FlashAttention, using a small block size can still hurt throughput due to CPU overheads, particularly due to the overhead of creating Block-Tables for every iteration (§3.3). For example, FlashAttention with block size 128 delivers 7% higher throughput than block size 16 for Llama-3-8B (531 vs 494 tokens per second with batch size 32).
@@
Figure 10. Latency of decode iterations with and without overlapping memory allocation with compute (batch size=4,context length=32K). Spikes show the latency impact of synchronous memory allocation.

### v2

@@ 5.4.2. Hiding memory allocation latency
The serving framework invokes the step API in every iteration. The latency of step depends on how many new pages need to be mapped in the virtual tensors of KV-cache. Consider, for example, that the KV-cache of one request needs to be extended for Yi-34B which has 60 layers. This requires 120 calls to cuMemMap + cuMemSetAccess each of which takes about 40 microseconds. Therefore, growing the KV-cache of one request by new pages (at all layers) adds about 5 millisecond latency to the corresponding iteration. Further, the latency overhead grows proportional to the number of requests that need new pages in a given iteration. We propose the following optimizations to hide the latency of allocation:
Overlapping memory allocation with compute
@@
Deferred reclamation + eager allocation
We observe that allocating physical memory for a new request can be avoided in many cases. Consider that a request R1 completed in iteration i and a new request R2 joins the running batch in iteration i+1. To avoid allocating new pages to R2 from scratch, vAttention simply defers the reclamation of R1’s pages and assigns R1’s reqId to R2. This way, R2 uses the same tensors for its KV-cache that R1 was using – which are already backed by physical pages (for security reasons, we still zero-fill such pages before allowing a new request to use them). Therefore, new pages are required only if the context length of R2 is bigger than that of R1.
We further optimize memory allocation by proactively allocating a small number of physical pages ahead of time. For this purpose, we try to keep a certain number of pages (configurable based on expected workload) mapped into the virtual tensors of one of the inactive reqId. When a new request arrives, we allocate this reqId and identify a new reqId to be allocated next and map physical pages for it. In most cases, these eager allocations obviate the need to allocate physical pages in the critical path of prefill execution. Finally, we trigger memory reclamation only when the number of physical pages cached in vAttention falls below a certain threshold (e.g., less than 10% of GPU memory). We delegate both deferred reclamation and eager allocation to the background thread that the step API spawns.
6. Evaluation
@@
Figure 6. Prefill throughput. vAttention backed systems outperform the paged counterparts of both FlashAttention and FlashInfer. Throughput for longer contexts is lower due to the quadratic complexity of prefill attention.
@@
  • How does vAttention perform for prefill and decode phases (§6.1, §6.2)?
  • How vAttention impacts LLM serving throughput (§6.3)?
  • What is the effect of each of our optimizations (§6.4)?
@@
Models and hardware: We evaluate three models Yi-6B (yi-6b-200k-hf, ), Llama-3-8B (llama-8b-hf, ) and Yi-34B (yi-34b-200k-hf, ), using a single NVIDIA A100 GPU for Yi-6B, and two NVLink-connected A100 GPUs for Llama-3-8B and Yi-34B (see Table 4). Each GPU has 80GB physical memory. We use tensor-parallelism degree of two (TP-2) for both Llama-3-8B and Yi-34B.
@@
Evaluation methodology: The computation and memory allocation pattern of the prefill and decode phases is substantially different (sarathiserve2024, ; splitfuse2024, ; distserve2024, ). Attention kernels used for these two phases are also different and hence we also evaluate them separately. The prefill phase requires one time memory allocation potentially spanning multiple pages. In comparison, the decode phase requires incremental memory allocation over the lifetime of a request (vllmsosp, ). We define throughput as the number of tokens processed (or generated) per second.
Serving framework: For a fair comparison, we use vLLM v0.2.7 as a common serving framework in all our experiments. We integrated state-of-the-art kernel libraries of both FlashAttention v2.5.9 (flashattention, ; flashattention2, ; fagithub, ) and FlashInfer v0.4.0 (figithub, ) as attention back-ends into vLLM, and further added support for dynamic memory allocation via vAttention to their non-paged kernels. While FlashAttention and FlashInfer are both based on the same underlying techniques (e.g., FlashDecoding (flashdecoding, )), they use different Block-Table formats; the former use a simple lookup table whereas the latter uses a compressed Block-Table to aid fetching KV-cache block locations.
Baselines: We compare performance obtained by using the non-paged attention kernels (backed by vAttention for dynamic memory allocation), and their paged counterparts. In addition, we also compare against vLLM’s decode kernel (note that vLLM does not have a paged prefill kernel). For a fair comparison, we also profiled each system to find its best performing configuration. Accordingly, we set KV-cache block size to 16 for both vLLM and FlashInfer, and 256 for FlashAttention. Using a higher block size for vLLM increases its kernel latency by up to 3×3$\times$3 × as shown in Figure 3, and using a smaller block size for FlashAttention paged kernel increases its latency by up to $9\%$ (Table 9). We find that the choice of page size does not affect vAttention as long as fragmentation is not a concern (Figure 11).
Model Context Length FA_Paged FA_vAttention FI_Paged FI_vAttention
Yi-6B 64K 10.6 (7.0) 9.1 (5.5) 10.9 (6.0) 9.1 (5.4)
128K 37.9 (30.3) 30.5 (23.1) 35.4 (25.4) 30.7 (23.3)
192K 81.5 (70.0) 64.6 (53.6) 73.0 (58.3) 65.1 (53.6)
@@
-8B 128K 20.4 (15.4) 16.8 (11.6) 20.3 (12.8) 16.8 (11.4)
192K 43.3 (35.6) 34.8 (26.9) 40.9 (29.7) 34.7 (26.7)
[…]
Yi-34B 128K 82.8 (56.9) 68.4 (43.2) 76.0 (46.7) 67.4 (42.5)
192K 170.7 (131.8) 136.9 (98.8) 148.8 (104.7) 134.6 (96.5)
Table 5. Prefill completion and attention (in parenthesis) time with different attention back-ends (unit: seconds).
Figure 7. Decode throughput. FA_vAttention is on par with FA_Paged (note the overlapping lines) which is the best among all PagedAttention based alternatives, while outperforming FI_Paged and vLLM.
6.1. Prefill Evaluation
We evaluate four configs for prefill: FA_Paged, FI_Paged, FA_vAttention and FI_vAttention. Configurations with the “_Paged” suffix represent the PagedAttention-based kernel and those with “_vAttention” use the non-paged kernels of the respective library. Figure 6 shows the prefill throughput for Yi-6B, Llama-3-8B and Yi-34B. We summarize our key findings below.
Small contexts: For small contexts, prefill cost is dominated by the linear operators i.e., attention’s contribution is relatively low (sarathi2023, ). Hence, even though vAttention helps speed up attention computation, the throughput of both paged and vAttention back-ends is nearly identical in FlashAttention. However, for FlashInfer, we find that using vAttention helps improve prefill throughput even for small contexts. This is because FlashInfer incurs various other sources of overhead in the paged version. First, appending a new K or V tensor to the KV-cache requires a single tensor copy operation in vAttention, whereas in a paged implementation, it requires appending one block at a time (the copy operation has been optimized for FlashAttention by vLLM (vllm-copy-kernels, )). Second, FlashInfer involves creation and deletion of a few objects for its compressed Block-Tables in every iteration. vAttention avoids such overheads because it maintains KV-cache’s virtual contiguity, eliminating the need for a Block-Table.
Long contexts: The contribution of attention computation becomes significant at 16K and higher context lengths in our experiments. Therefore, even for FlashAttention back-end, vAttention outperforms the paged counterpart. For example, at context length 192K, FA_vAttention outperforms FA_Paged by 1.24×1.24$\times$1.24 ×, 1.26×1.26$\times$1.26 × and 1.24×1.24$\times$1.24 × for Yi-6B, Llama-3-8B and Yi-34B, respectively. Similarly, for FlashInfer back-ends, FI_vAttention improves prefill throughput by up to 1.25×1.25$\times$1.25 × and 1.36×1.36$\times$1.36 × for Yi-6B and Llama-3-8B (context length 16K), and 1.17×1.17$\times$1.17 × for Yi-34B (context length 32K).
Attention time: vAttention’s improvement in prefill throughput is primarily due to faster attention kernels enabled by a (virtually) contiguous KV-cache. This is because the prefill phase of a long prompt is primarily dominated by attention computation, as can be observed by comparing the numbers inside parenthesis with the total prefill completion time in Table 5. For FlashAttention, nearly all the gains of vAttention are due to faster attention kernels e.g., prefill gains of 1.5 seconds, 7.4 seconds and 16.9 seconds for Yi-6B are almost entirely due to gains in attention computation. vAttention enabled kernels also help with the FlashInfer back-end. In addition, FI_Paged also has other sources of overheads e.g., in the 14 seconds of total savings (Yi-34B, 192K context), only 7 seconds is due to attention and the rest is due to other sources.
Figure 8. Makespan of serving a set of 50 requests with varying context lengths and the ratio of prefill to decode tokens.
6.2. Decode Evaluation
For decodes, in addition to FA_Paged and FI_Paged, we also evaluate the throughput obtained with vLLM’s decode kernel (the first ever kernel to support PagedAttention). For vAttention, we use FlashAttention’s non-paged kernel. Unfortunately FlashInfer’s non-paged decode kernel has significantly higher latency compared to all the other kernels we used in these experiments (up to 14.6×14.6$\times$14.6 ×, Table 10). Hence, while vAttention supports dynamic memory allocation for FlashInfer’s non-paged decode kernel, we omit it for evaluation in this section (note that vAttention is responsible only for allocating memory – it cannot make a slow kernel fast).
@@ 6.2. Decode Evaluation
Figure 7 shows the decode throughput of Yi-6B, Llama-3-8B and Yi-34B with varying batch size up to 32 (except for Yi-34B which runs out of memory for batch size 32). We set the initial context length of each request to 16K tokens and calculate decode throughput based on the mean latency of 400 decode iterations (see Figure 7).
First, vAttention is on par with the best of PagedAttention as shown by FA_Paged and FA_vAttention in Figure 7. In comparison, FI_Paged has somewhat lower throughput and vLLM is the worst for all models and configurations. For example, FA_Paged and FA_vAttention outperform vLLM by up to 1.99×1.$9$, 1.58×1.58$\times$1.58 × and 1.53×1.53$\times$1.53 × for Yi-6B, Llama-3-8B and Yi-34B, respectively. The primary reason is that vLLM’s decode kernel has significantly higher latency than the FlashAttention based kernels; while FlashAttention has continuously adopted new optimizations (e.g., FlashDecoding (flashdecoding, )), vLLM has lagged behind. For example, Table 6 shows that vLLM’s PagedAttention kernel incurs up to 2.8×2.8$\times$2.8 ×, 1.5×1.5$\times$1.5 ×, and 2.5×2.5$\times$2.5 × higher latency than FlashAttention kernels for Yi-6B, Llama-3-8B and Yi-34B. This is despite vLLM being in an actively maintained open-source serving stack, as well as being used by various companies for serving LLMs. This is an important result that underlines the importance of adopting new optimizations.
Model BS vLLM FA_Paged FI_Paged FA_vAttention
Yi-6B 16 32.3 11.5 15.2 11.3
32 64.1 25.5 25.4 25.3
[…]
32 35.3 25.4 23.23 25.3
[…]
16 55.1 21.7 28.8 21.8
Table 6. Total latency of attention kernel (sum of all layers) per decode iteration (in milliseconds, BS = batch size).
Second, relative gains of FA_vAttention and FA_Paged increase over vLLM with the batch size e.g., as batch size increases from 4 to 32 for Llama-3-8B, relative gains increase from 1.05×1.05$\times$1.05 × to 1.58×1.58$\times$1.58 ×. This is because the latency of a decode attention kernel is proportional to the total number of tokens in the batch (vidur, ). Therefore, the contribution of attention kernel in the overall latency – and consequently gains with a more efficient kernel – increase with the batch size. Further, for the same reasons as discussed in §6.1 (faster attention and lower CPU overhead), FA_vAttention delivers up to 1.23×1.23$\times$1.23 × higher throughput than FI_Paged (Yi-6B, batch size 12).
Finally, note that vAttention is only as good as the state-of-the-art PagedAttention for decode throughput, as compared to prefills where it outperforms PagedAttention. We believe this is due to memory bound nature of decode attention i.e., memory stalls make it possible to hide the effect of additional compute that paging support requires. This is evident from Table 6 wherein FA_Paged and FA_vAttention have similar latency. However, hiding compute overhead in a prefill attention kernel is hard because it is already compute bound (see Table 5 for latency gains in the prefill attention kernel).
6.3. End-to-end Throughput Evaluation
For end-to-end performance, we measure the makespan of serving a static trace of 50 requests for long context scenarios. To understand system performance for different workloads, we vary the initial context length of requests from 32K to 128K, and the ratio of prefill to decode tokens (referred as P:D) from 500 to 50. Note that a higher P:D as well as a longer context indicates that the workload is more prefill bound; we expect vAttention to improve performance in such cases.
Figure 8 shows the makespan for all the three models. For Yi-6B, FA_vAttention is up to 1.22×1.$2$, 1.16×1.16$\times$1.16 × and 1.13×1.13$\times$1.13 × faster than FA_Paged for P:D ratio of 500, 100 and 50. FA_vAttention is also up to 1.19×1.19$\times$1.19 × faster for Llama-3-8B and up to 1.16×1.16$\times$1.16 × faster for Yi-34B, compared to FA_Paged. Our gains over FA_Paged are higher at higher P:D ratios, and for a fixed P:D, our gains increase as the context length grows. These results are consistent with our expectation since a higher P:D and context lengths both mean that the system spends more time processing prefills where use of non-paged attention kernels outperform paged kernels.
Interestingly, we observe that despite being faster than FA_Paged in prefills, FI_Paged under-performs FA_Paged in end-to-end throughput in most cases (due to inefficient decodes as shown in Figure 7). Hence, our gains over FI_Paged are relatively higher (up to 1.29×1.29$\times$1.29 ×), except for some cases where FI_Paged outperforms FA_Paged (e.g., Yi-6B and Yi-34B, context length=128K, P:D=500).
@@ 6.3. End-to-end Throughput Evaluation
Figure 9. Latency of decode iterations with and without overlapping memory allocation with compute (batch size=32, context length=4K–8K, model: Llama-3-8B)
6.4. Ablation Studies
6.4.1. Hiding allocation latency
Figure 9 shows the latency of more than 2500 decode iterations of Llama-3-8B. For this experiment, we used a batch of 32 requests and initialized the prefill context length of each request to be in between 4K–8K (chosen randomly). It is evident that overlapping memory allocation with compute effectively hides the latency of allocating memory via CUDA APIs. We used 2MB pages for this experiment to show that even the worst case latency can be hidden (Table 2 shows that 2MB pages incur highest latency). In contrast, allocating memory synchronously via CUDA APIs leads to frequent latency spikes of 5ms–20ms, depending on how many requests need new physical memory pages.
6.4.2. Deferred reclamation
Figure 10 shows that synchronous memory allocation using CUDA APIs for prefills incurs overhead of up to 1.15×1.15$\times$1.15 × with 64KB pages and up to 1.03×1.03$\times$1.03 × with 2MB pages. In most cases, deferred reclamation eliminates the need to invoke CUDA APIs for prefills because a newly arrived request can simply re-use physical memory pages that were allocated to a prior request. This way, deferred reclamation ensures that prefill latency is not affected by memory allocation.
6.4.3. Memory allocation bandwidth
Table 7 shows that even with 64KB (our smallest) page size, vAttention can allocate as much as 7.6GB per second per GPU. This is more than an order of magnitude higher than the maximum memory allocation rate of 600MB per second of decodes (Figure 4). Larger page sizes and higher TP dimensions increase the memory allocation rate proportionally. Therefore, memory allocation bandwidth of CUDA APIs is more than enough for LLM inference.
Figure 10. Prefill completion time of a single prompt of 16K tokens with different memory allocation strategies.